Wednesday, 26 August

Wednesday, 26 August

Invited Lectures

\underline{14:00 – 14:45}

01 Stevo B. Todorcevic, CNRS, Université de Paris 7, Paris, France, Mathematicki Institut, Beograd, Yugoslavia, and University of Toronto, Toronto, Canada Basis Problems in Combinatorial Set Theory (I: p.35) Chair: Martin Goldstern

MA 005

05 Dmitri Burago, Pennsylvania State University, University Park, USA Hard Balls Gas and Alexandrov Spaces of Curvature Bounded Above (I: p.68) Chair: Frank Duzaar

MA 001

11 Antonio Giorgilli, Università degli Studi di Milano, Milano, Italy On the problem of stability for near to integrable Hamiltonian systems (I: p.132) Chair: Johannes Kellendonk

H 1028

13 Alain Lascoux, CNRS, Université de Marne La Vallée, Noisy-le-Grand, France Ordering the symmetric group: Why use Iwahori-Hecke algebras? (I: p.157) Chair: Andrei Zelevinsky

H 2032

15 Lloyd N. Trefethen, Oxford University, Oxford, UK Schwarz–Christoffel Mapping in the Computer Era (I: p.181) Chair: Peter J. Deuflhard

H 1012

18 Michèle Artigue (speaker), Université de Paris 7, Paris, France De la compréhension des processus d’appretissage à la conception de processus d’enseighement (I: p.202) Hans-Georg Steiner (respondent), Universität Bielefeld, Bielefeld, Germany “From the Understanding of Learning Processes to the Design of Teaching Processes” (I: p.203) Chair: Hans-Christian Reichel

H 104

\underline{15:00 – 15:45}

03 Takeshi Tsuji, RIMS, Kyoto University, Kyoto, Japan p-adic Hodge Theory in the semi-stable reduction case (I: p.55) Chair: Helmut Koch

MA 005

05 Gerhard Huisken, Universität Tübingen, Tübingen, Germany Evolving hypersurfaces by their curvature (I: p.72) Chair: Frank Duzaar

MA 001

12 David Oliver Siegmund, Stanford University, Stanford, USA Genetic linkage analysis: an irregular statistical problem (I: p.148) Chair: Uwe Küchler

H 1028

13 Andrei V. Zelevinsky, Northeastern University, Boston, USA Multisegment duality, canonical bases and total positivity (I: p.163) Chair: Mikhail M. Kapranov

H 2032

15 Bjorn Erik Engquist, UCLA, Los Angeles, USA Wavelet Based Numerical Homogenization (I: p.179) Chair: Ralf Kornhuber

H 1012

\underline{16:15 – 17:00}

01 Gregory Hjorth, UCLA, Los Angeles, USA When is an equivalence relation classifiable? (I: p.32) Chair: Ronald B. Jensen H 1012 03 Shou-Wu Zhang, Columbia University, New York City, USA Small points, Heegner points, and Arakelov theory (I: p.57) Chair: Jürg Kramer

MA 005

08 Pertti Mattila, University of Jyväskylä, Jyväskylä, Finland Rectifiability, analytic capacity, and singular integrals (I: p.100) Chair: Srishti (S.D.) Chatterji

MA 001

12 Boris Tsirelson, Tel Aviv University, Tel Aviv, Israel Within and beyond the reach of Brownian innovation (I: p.149) Chair: Yuri Kifer

H 1028

16 Frank Hoppensteadt, Arizona State University, Tempe, USA Canonical Models in Mathematical Neuroscience (I: p.187) Chair: Karl Sigmund

H 2032

\underline{16:15 – 18:00}

18 Miguel de Guzman, Universidad Complutense de Madrid, Madrid, Spain Bernard R. Hodgson, Université Laval, Québec, Canada Aline Robert, IUFM de Versailles, France Vinicio Villani (moderator), Università di Pisa, Pisa, Italy Round table on: Difficulties in the Passage from Secondary to Tertiary Education (I: p.204) Chair:

H 104

\underline{17:15 – 18:00}

01 Alex James Wilkie, University of Oxford, Oxford, England O-minimality (I: p.35) Chair: Anand Pillay

H 1012

03 Hans Peter Schlickewei, Universität Marburg, Marburg, Germany The Absolute Subspace Theorem and Applications (I: p.54) Chair: Alfred J. van der Poorten

MA 005

08 Stefan Müller, Max-Planck-Inst. für Math. in den Naturwissenschaften, Leipzig, Germany Unexpected solutions of first and second order partial differential equations (I: p.104) Chair: John M. Ball

MA 001

12 Iain Murray Johnstone, Stanford University, Stanford, USA Oracle Inequalities and Nonparametric Function Estimation (I: p.146) Chair: Persi W. Diaconis

H 1028

16 Charles S. Peskin, Courant Institute, New York University, New York, USA How to Catch and Transport a Chromosome (I: p.188) Chair: Willi Jäger H 2032